40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  Cryo Lab eLog  Not logged in ELOG logo
Entry  Thu Jan 28 18:03:53 2021, shruti, DailyProgress, PSOMA, cavity locking Path.pdf
    Reply  Mon Feb 1 18:25:47 2021, shruti, DailyProgress, PSOMA, cavity locking 
       Reply  Wed Feb 3 10:56:54 2021, shruti, Update, PSOMA, cavity locking 
          Reply  Thu Feb 4 16:19:35 2021, shruti, Update, PSOMA, cavity alignment 5FC37CEF-5B30-4FF0-ADF6-9FD4C59A1027.jpeg
             Reply  Thu Feb 4 22:27:43 2021, rana, Update, PSOMA, cavity alignment IMC_HOManalysis.ipynb
                Reply  Tue Feb 9 11:14:48 2021, shruti, Update, PSOMA, cavity alignment 
                   Reply  Tue Feb 16 11:30:14 2021, shruti, Update, PSOMA, Beat recovered, fiber organization, alignment F2E33DA5-DB03-46EC-B341-E388D96F1BC5.jpeg069D0D53-9111-4B33-96AD-05B907C4BBCB.jpegBeat_low_freq_oscillations.mp4
Message ID: 2646     Entry time: Tue Feb 16 11:30:14 2021     In reply to: 2632
Author: shruti 
Type: Update 
Category: PSOMA 
Subject: Beat recovered, fiber organization, alignment 

I played around with the temperature and current to recover the beat (Attachment 3):

south laser temp: 9.373 kOhm, north laser temp: 9.602 kOhm

The low frequency oscillation still remains.

 

I decided to re-organize the fiber to make the north and south paths more evident visually. In the process, probably because I could not quite see the fiber coupler and was inserting it incorrectly or I somehow damaged the fiber coupler, the north coupler on the ADAFCB4 did not mode-match adequately so I replaced it with a single coupler borrowed from the west path in Johannes' setup. Now both paths at the output of the fiber show ~2 mW power and the beat is at around -15 dBm. (Attachment 1).

 

I've replaced the south path fiber to free-space coupler mount and post with an adjustable one to get the beam height to 4 inches. (Attachment 2)

 

 

Attachment 1: F2E33DA5-DB03-46EC-B341-E388D96F1BC5.jpeg  3.566 MB  | Hide | Hide all
F2E33DA5-DB03-46EC-B341-E388D96F1BC5.jpeg
Attachment 2: 069D0D53-9111-4B33-96AD-05B907C4BBCB.jpeg  3.546 MB  | Hide | Hide all
069D0D53-9111-4B33-96AD-05B907C4BBCB.jpeg
Attachment 3: Beat_low_freq_oscillations.mp4  4.086 MB
ELOG V3.1.3-